# Causal Discovery in Financial Markets: A Framework for Nonstationary Time-Series Data

### v2

@@ 3.2 Testing the Assumptions
On the other hand, for the linearity assumption, we modify an algorithm used in prior work (Peters et al., 2014) used in order to test post-hoc if the relationships are linear. Specifically, the relationships are assumed to be linear in conjunction assuming with additive noise. The algorithm used in Peters et al. (2014) is to perform a regression and then an independence test on the noise. Say the null hypothesis we are testing is that $X$ and $Y$ have a linear relationship given $Z$ (where we allow for $Z$ to have a non-linear relationship with $Y$ and the noise to be heteroscedastic with respect to $Z$), we perform a Gaussian process regression $\hat{f}$, where we use the linear kernel for $X$ and a non-linear kernel (e.g., RBF) for $Z$. Having done that, we then perform a CI test on $Y-\hat{f}(X,Y)\perp\!\!\!\perp X|Z$. Note that our test is more generic than testing if all relationships are linear.
@@ 3.3 Importance of Modeling Lagged Relationships
$P(x_{t}|x_{u})$ $=P(x_{t}|\ \mathcal{N}(x_{u}e^{-\mathinner{\!\left\lvert t-u\right\rvert}},e^{-\mathinner{\!\left\lvert t-u\right\rvert}}))\,,\,u<t\,.$
@@ 3.4 Comparison to Algorithm in CD-NOD
Figure 3: Diagram of example causal graph for which CD-NOD’s time-series algorithm would make a mistake.
@@ 4 Evaluation
Figure 4: F-score evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
@@
In our algorithm, one major design decision is the conditional independence test. In order to ascertain which conditional independence test might be most effective, we test our algorithm on different simulated datasets. In total, we tested graphs with 3, 4, 5, 6, 8, 10, and 15 nodes, for a total of 350 graphs (50 graphs per number of nodes). For each graph, we tested having 50, 150, 300, 500, and 1,000 data points.
For the conditional independence tests, we specifically test using partial correlations (ParCorr, a conditional independence test that assumes linearity), Kernel-based Conditional Independence Test (KCIT), Randomized Conditional Correlation Test (RCoT), and Conditional Mutual Information K-nearest neighbor (CMIknn) Test. For both KCIT and RCoT, there are a few variations in how the p-value can be computed, specifically using the Satterthwaite–Welch (SW) method (Welch, 1938; Satterthwaite, 1946) or Hall– Buckley–Eagleson (HBE) method (Hall, 1983; Buckley and Eagleson, 1988). The goal of these algorithms is to approximate the CDF of a weighted sum of squared normals, which is the limiting behavior of the test statistic for KCIT and RCoT. The hyperparameters used for each test can be found in Appendix A.
In Figure 4, we compare the average F-score for each CI test across a varying number of nodes in the graph and a varying number of data points per graph. As expected, the F-score improves as the number of data points increase. Possibly most interesting is that ParCorr performs the best for low data regimes. Further, we can see, as found by previous work (Bodenham and Adams, 2015), that the SW variant of KCIT tends to be weaker than HBE (except when there are a large number of nodes and only 50 data points). However, there seems to be little difference for RCoT. So, relying on results in Bodenham and Adams (2015), we recommend the HBE variant. Finally, while CMIknn is quite competitive, we do not recommend its usage due to its lengthy runtime (Figure 5).
@@ 5 Case Studies
We apply CD-NOTS on three different datasets: 1) Fama-French factors and Apple’s returns, 2) unemployment, CPI, and PPI of different countries, and 3) Price-to-Book ratio and stock returns of financial companies in the S&P 500. Most of the work done in factor investing and economic indices’ analysis rely on the correlation between the variables and do not consider the fundamental question of “why” (López de Prado, 2023). In each case study, we explain the data and properties we see in the data or has been found in prior work, discuss the implications of the causal networks found by CD-NOTS, and, finally, show that the assumptions of stationarity, linearity, and no lagged relationships would be incorrect to assume and thus, any algorithm making said assumptions would be inaccurate and invalid to use on these datasets.
For all the graphs in the section, we color by the maximum p-value for that edge across all tests run during the causal skeleton step (Stage 2). While we recognize that p-values are not a metric to measure the strength of the relationship, we can think of the vicinity of the p-value to 0.05 as a measure of the sensitivity of the graph to that hyperparameter (i.e., the threshold).
@@ 5.1 Fama-French Factors
Figure 7: Time-series of Fama-French factors and Apple’s returns.
Refer to caption Refer to caption Refer to caption Refer to caption (a) Jan 2000 - Dec 2007 Refer to caption (b) Jan 2008 - Dec 2015 Refer to caption (c) Jan 2016 - Dec 2022
Figure 8: Results of CD-NOTS KCIT run on three non-overlapping time periods.
Refer to caption Refer to caption
Figure 9: Results of CD-NOTS KCIT run from beginning of 2000 to end of 2022.
@@
For our first case study, we analyzed the relationship between the Fama-French factors (Fama and French, 1993) and Apple’s returns from the beginning of 2000 to the end of 2022. In the data in Figure 7, we show that there is some correlation between the variables (e.g., the spike in volatility around 2008 can be seen in all six subplots) and, accordingly, all variables show volatility clustering.
@@
In Figure 8, we ran CD-NOTS using KCIT (RCoT results are in Appendix B) on three non-overlapping periods of six to seven years (approximately 2000 observations each). We note in all three periods that there is no connection from $T$ to Apple’s returns (AAPL_RET), implying that the nonstationarity observed in Apple’s returns can be explained away by the Fama-French factors. Further, the lagged relationships tends to be sparse, with many of the found lagged relationships having a high p-value (close to 0.05). However, we see that the relationship between the factors seems to change over time, as well as the relationship between the factors and Apple’s returns. In Figure 9, we run CD-NOTS with KCIT on the whole time range, and we can see now that that there is a connection from $T$ to Apple’s returns (AAPL_RET). This is to be expected for two reasons: 1) over the course of the last 22 years, Apple’s business model has changed, leading to exposures to different factors over time, and 2) the changing relationship between the factors and Apple’s returns, as seen in Figure 8.
@@
One of the assumptions of CD-NOTS is causal consistency that the causal graph does not change over time. We assert that the change in dependencies between the factors and Apple’s returns that we see in Figure 8 are not a result of changes in the causal directions but in the strength of the causal relationships as captured by the nonstationarity. Further, we note that the fact that there is a relationship between $T$ and all the Fama-French factors (nonstationarity) implies that any causal discovery algorithm assuming stationarity would give invalid results. Similarly, we find statistically significant lagged relationships. Further, for our linearity test, we test the linearity of Apple’s returns and each parent factor (CMA and Mkt-RF) and find the linearity hypothesis can be rejected for both factors; in other words, assuming linearity in the relationships between variables would be incorrect and thus could lead to incorrectly identified causal relationships.
@@ 5.2 Economic Data
Refer to caption (a) U.S. Refer to caption (b) Canada Refer to caption (c) Japan Refer to caption (d) France Refer to caption (e) United Kingdom Refer to caption (f) Italy
Figure 10: Results of CD-NOTS ParCorr run on six countries’ economic variables.
Figure 11: Results of CD-NOTS KCIT run on the economic variables of multiple countries.
@@
For our second case study, we analyzed the relationship between the month-over-month percentage change in CPI, the month-over-month percentage change in PPI, and unemployment for the U.S., Japan, Canada, India, Italy, the U.K., and France. We use monthly data from the beginning of 2000 to the end of 2023.
@@
In Figure 10, we run CD-NOTS with ParCorr on the U.S., Japan, Canada, Italy, France, and the U.K.; we use ParCorr since we have no more than 300 observations per country. We can see that, while the graphs are not identical, they are quite similar, especially if we ignore edges with a higher p-value. Further, we notice the three graphs for France, the U.K., and Italy are very similar, possibly due to their geographic and economic proximity. We repeat the experiment with all countries using KCIT (Figure 11); further, we introduce another static variable (Country) into the graph. Similar to the previous case study in which the graph over all time periods is akin to a superset of the edges of individual time periods, we see a similar behavior here. Note that Country does not point to any other variable suggesting that the behavior across countries is independent of the country; this observation could be used to justify training a single econometric model across countries.
@@
Similar to the previous case study, we see statistically significant lagged relationships and statistically significant nonstationarity, implying we would not be able to trust the results of an algorithm that cannot handle these types of properties. For our linearity test, we test the linearity of the percentage change of CPI and each parent node (both including and excluding the contemporaneous percentage change in PPI since its direction was not inferred). We find that the only time the linearity hypothesis cannot be rejected is when using the contemporaneous percentage change in PPI as a parent and checking the linearity with the one-lag percentage change of CPI. We note that this is inconsistent with our usage of ParCorr in Figure 10 since ParCorr is a linear-based CI test. In Appendix C, we show the results of using KCIT and RCoT where we find qualitatively that the graphs are similar.
@@ 5.3 Company Financials and Returns
Refer to caption (a) Contemporaneous Effects Refer to caption (b) Lagged Effects Refer to caption (c) Contemporaneous Effects Refer to caption (d) Lagged Effects Refer to caption (e) Comparison of Returns and N. LPTB Refer to caption (f) Cumulative Returns of Strategies
Figure 12: Results of CD-NOTS KCIT run on financials data for log Price-to-Book ratio (LPTB) and stock returns (Ret). Normalized LPTB (N. LPTB) refers to normalizing the LPTB by the mean and standard deviation of the LPTB of financial companies in that quarter. For Figure 11(f), our strategy is long-ing financial companies that have a normalized LPTB greater than the threshold which we compare against the market-cap weighted portfolio.
@@
For our third case study, we analyzed the relationship between the Price-to-Book ratio and returns of financial companies in the S&P 500. We use data from the beginning of 2010 to the end of 2023. We utilize reported quarterly numbers, where we set the date to be the end of each quarter. To ensure no look-ahead bias when defining the Price-to-Book ratio, for any given date (i.e., end of quarter), we use the Price-to-Book ratio available that day. We join the next quarter returns with this company financials information. For our experiments, we use the log Price-to-Book ratio (LPTB) as well as the normalized LPTB, which we define as the LPTB normalized by the mean and standard deviation of the LPTB of financial companies in that quarter.
@@
In Figure 11(b) and Figure 11(d), we compare the graphs we get running CD-NOTS using KCIT using normalized LPTB instead of LPTB. We find that the nonstationarity is removed (T does not point to N. LPTB). Further, we see that the previous quarter’s normalized LPTB point to the next quarter’s returns with a reasonable p-value suggesting that its inclusion in the graph is not purely a function of the threshold we chose. In Figure 11(e), we compare the next quarter returns and the normalized LPTB and see that the there is a small trend in the mean (the black circles). To this end, we create a simple trading strategy where we invest in the subset of the financial securities whose LPTB is greater than some threshold. In Figure 11(f), we see this simple approach outperforms a market-cap weighted strategy; prior work (Monge et al., 2023) has found that growth stocks have been outperforming value stocks. Importantly, we note that the goal of this is to not introduce a new signal but to show how causal discovery can be a tool in searching for statistically significant relations which, if the assumptions are true for the data, can be interpreted as causal.
@@
Similar to previous case studies, we see statistically significant lagged relationships and statistically significant nonstationarity, implying we would not be able to trust the results of an algorithm that cannot handle these types of properties. For our linearity test, we test the linearity of the normalized LPTB and returns conditional on all other parent nodes of returns and find the p-value be to 0.0 (i.e., we can reject the null hypothesis that the relationship is linear).

### v3

@@ 2.2.1 Objective Assumptions
While many recent algorithms do not assume linearity, for comprehensive analysis, we test this assumption post-hoc in our case studies (Section 4) using a modified version of a previously employed algorithm (Peters et al., 2014). Typically, prior work assuming linear relationships also assume additive noise. Accordingly, when we test for linearity, we also test for additive noise. The algorithm used in Peters et al. (2014) involves performing a regression followed by an independence test on the noise. For instance, if the null hypothesis we are testing is that $X$ and $Y$ have a linear relationship given $Z$ (where we allow for $Z$ to have a non-linear relationship with $Y$ and the noise to be heteroscedastic with respect to $Z$), we perform a Gaussian process regression $\hat{f}$ using a linear kernel for $X$ and a non-linear kernel (e.g., RBF) for $Z$. Subsequently, we perform a CI test on $Y-\hat{f}(X,Y)\perp\!\!\!\perp X|Z$. Note that our test is more generic than testing if all relationships are linear.
2.2.2 Subjective Assumptions
Almost every causal discovery algorithm is built upon subjective assumptions. It is crucial to justify these assumptions due to their subjective nature. Here, we provide an intuitive explanation of these assumptions in order to allow readers to judge the realism of these assumptions.
Pseudo-sufficiency
Any potential confounder described as a smooth function of time can be captured by the time-indexed node; while, if there exists a confounder which is stationary and not changing through time, the algorithm will not capture it.
Consistency
The consistency assumption states that the causal links remain unchanged through time. We assert any deviations observed in the causal graph over time in Section 4 can be attributed to sampling variations and data availability.
Faithfulness
This assumption can be violated in many ways, including xor connections (or gating mechanisms), deterministic functions, or cancelling paths (Marx et al., 2021). In our case studies, we do not have situations of deterministic functions which can be directly tested from the data.
Randomness
The randomness assumption comes from the CI tests which enables the derivation of the null distribution (statistical validation). We note that this assumption does not hold for time series [^3] This is a limitation for all constraint-based causal discovery algorithms that use p-values of CI tests to define a threshold, and this is a gap to be filled in future work.; this implies that the p-value threshold used does not accurately control the false positive rate. We argue that testing the robustness of our algorithm to different p-value thresholds mitigates this concern.
@@ 2.3 Importance of Modeling Lagged Relationships
[…]
@@ 2.4 Comparison to Algorithm in CD-NOD
Figure 3: Diagram of example causal graph for which CD-NOD’s time series algorithm would make a mistake.
@@ 3 Evaluation
One major design decision in any constrained-based algorithms, including ours, is the CI test. In order to ascertain which test might be most effective, we test our algorithm on different simulated datasets. In total, we tested graphs with 3, 4, 5, 6, 8, 10, and 15 nodes, for a total of 350 randomly generated graphs (50 graphs per number of nodes). For each graph, we tested having 50, 150, 300, 500, and 1,000 data points. The relationship between any connected nodes can be either linear or non-linear (quadratic, exponential, or sine). The max lag we allowed was five.
For the conditional independence tests, we specifically test using partial correlations (ParCorr, a conditional independence test that assumes linearity), KCIT, RCoT , and CMIknn. For both KCIT and RCoT, there are a few variations in how the p-value can be computed, specifically using the Satterthwaite–Welch (SW) method (Welch, 1938; Satterthwaite, 1946) or Hall–Buckley–Eagleson (HBE) method (Hall, 1983; Buckley and Eagleson, 1988). The goal of these algorithms is to approximate the CDF of a weighted sum of squared normals, which is the limiting behavior of the test statistic for KCIT and RCoT. The hyperparameters used for each test can be found in Appendix A.
In Figure 4, we compare the average F-score [^4] F-score is the harmonic mean of precision and recall metrics i.e., F-score$=2\,\frac{precision\times recall}{precision+recall}\,$. for each CI test across a varying number of nodes in the graph and a varying number of data points per graph. As expected, the F-score improves as the number of data points increase. Possibly most interesting is that ParCorr performs the best for low data regimes. Further, we can see, as found by previous work (Bodenham and Adams, 2015), that the SW variant of KCIT tends to be weaker than HBE (except when there is a large number of nodes and only 50 data points). However, there seems to be little difference for RCoT. Hence, relying on results in Bodenham and Adams (2015), we recommend the HBE variant.
Figure 4: F-score evaluation for CD-NOTS with different CI tests, tested out on different simulated datasets.
From our simulation studies, we recommend using RCoT for large datasets since it is much faster than KCIT on a CPU. However, we do note that the difference in speed is much less when using a GPU implementation of KCIT and RCoT; thus, if a GPU is available, we recommend KCIT since it does not have a randomized component. Finally, while CMIknn is quite competitive, we do not recommend its usage on large problems due to its lengthy runtime.
Another important metric to consider while choosing a CI test is the computational cost of it. Figure 5 shows the average runtime of different tests for different sample sizes. In general, methods like CMIKnn are several orders of magnitude slower than simple linear tests like ParCorr which limits their applicability in very large samples. In terms of computations, KCIT is positioned in the middle and its cost grows in $\mathcal{O}(n^{3})$ with sample size. RCoT has time complexity of $\mathcal{O}(m^{2}n)$ where $m$ is the number of Fourier terms (typically in order of 10-100) set by the user which gives it a considerable computation saving when compared to KCIT. In general, as expected the number of conditions that needs to be tested grows rapidly with the number of nodes as shown in Figure 6. Note that the number of tests needed for a given network size will not be exactly the same for different CI tests due to fact that any edge removal in early states of skeleton discovery (which depends on the p-value generated by the given test) will result in a substantial decrease in condition sets that need to be tested and hence the total number of tests.
@@
Figure 6: Number of tests run for CD-NOTS for different CI tests. Note the y-axis is in log scale.
Another important parameter of the algorithm is the significance level $\alpha$. We ran a series of experiments with $\alpha$ levels [0.01, 0.05, 0.1] to study the sensitivity of the algorithm to the choice of $\alpha$. The results for F-score have been shown in Figure 7. Please see Appendix D.2 for other metrics. As expected, a lower $\alpha$ results in higher precision across all data ranges and CI tests. The only exception to this is RCoT in low data regimes, but this is expected because RCoT, by design, assumes there are large number of data points and its results are not reliable for small datasets. In terms of F-score, larger $\alpha$ values perform better in low data regimes; however, in large data regimes, small alphas perform better.
(a) ParCorr
(b) CMIKnn
(c) KCIT
(d) RCoT
Figure 7: F-score evaluation, comparing different confidence levels.
@@ 4 Case Studies
Most of the work done in factor investing and economic indices’ analysis rely on the correlation between the variables and do not consider the fundamental question of “why” (López de Prado, 2023). We apply CD-NOTS to three financial datasets: 1) Fama-French factors (Fama and French, 1993) and Apple’s returns, 2) unemployment, CPI, and PPI of different countries, and 3) Price-to-Book ratio and stock returns of financial companies in the S&P 500. In each case study, we explain the data and properties we see in the data or that has been found in prior work, discuss the implications of the causal networks found by CD-NOTS and, finally, show that the assumptions of stationarity, linearity, and no lagged relationships would be incorrect to assume and thus, any algorithm making said assumptions would be inaccurate to use on these datasets.
For all the graphs in this section, we color by the maximum p-value for that edge across all tests ran during the causal skeleton step (Stage 2). While we recognize that p-values are not a metric to measure the strength of the relationship, we can think of the vicinity of the p-value to 0.05 as a measure of the sensitivity of the graph to that hyperparameter (i.e., the threshold). While in our experiments in Section 3 we found that using a p-value of 0.01 to be better for larger datasets, we run our experiments with a p-value of 0.05, both to be consistent with prior work (Strobl et al., 2017) as well as to show the edges that would be removed by changing from 0.05 to 0.01.
@@ 4.1 Fama-French Factors
For our first case study, we analyze the relationship between the Fama-French factors and daily Apple’s returns from the beginning of 2000 to the end of 2022. In the data in Figure 9, we show that there is some correlation between the variables (e.g., the spike in volatility around 2008 and during COVID Pandemic can be seen in all six subplots) and, accordingly, all variables show volatility clustering.
Figure 9: Time series of Fama-French factors and Apple’s returns.
@@
In Figure 10 and Figure 11, we run CD-NOTS with KCIT on the whole time range with and without the $T$ node, respectively, or, in other words, test the impact of modeling nonstationarity. While the number of edges in the contemporaneous graph is about the same (one edge difference, specifically between SMB and CMA), there are fewer lagged relationships in Figure 10. This implies the dependence we see through time is better explained as a nonstationary relationship.
It should be noted that there is a connection from $T$ to Apple’s returns (AAPL_RET). This is to be expected for two reasons: 1) over the course of the last 22 years, Apple’s business model has changed, leading to exposures to different factors over time, and 2) the changing relationship between the factors and Apple’s returns.
Figure 10: Results of CD-NOTS KCIT run from beginning of 2000 to end of 2022. Left Contemporaneous graph Right Full time series graph
Figure 11: Results of CD-NOTS KCIT run from beginning of 2000 to end of 2022 without the time-indexed node. Left Contemporaneous graph Right Full time series graph
@@
We note that the fact that there is a relationship between $T$ and all the Fama-French factors (nonstationarity) implies that any causal discovery algorithm assuming stationarity would give invalid results. Similarly, we find statistically significant lagged relationships. Further, for our linearity test, we test the linearity of Apple’s returns and each parent factor (CMA and Mkt-RF) and find the linearity hypothesis can be rejected for both factors; in other words, assuming linearity in the relationships between variables would be incorrect and thus could lead to incorrectly identified causal relationships.
@@ 4.2 Economic Data
For our second case study, we analyze the relationship between the month-over-month percentage change in Consumer Price Index (CPI), the month-over-month percentage change in Producer Price Index (PPI), and unemployment for the U.S., Japan, Canada, India, Italy, the U.K., and France. We use monthly data from the beginning of 2000 to the end of 2023.
(a) U.S.
(b) Canada
(c) Japan
(d) France
(e) United Kingdom
(f) Italy
Figure 12: Results of CD-NOTS ParCorr ran on six countries’ economic variables.
@@
In Figure 12, we run CD-NOTS with ParCorr on the U.S., Japan, Canada, Italy, France, and the U.K.; we use ParCorr since we have no more than 300 observations per country. We can see that, while the graphs are not identical, they are quite similar, especially if we ignore edges with a higher p-value. Further, we notice the three graphs for France, the U.K., and Italy are very similar, possibly due to their geographic and economic proximity. We repeat the experiment with all countries using KCIT (Figure 13); further, we introduce another static variable (Country) into the graph. Similar to the previous case study in which the graph over all time periods is akin to a superset of the edges of individual time periods, we see a similar behavior here. One difference we see is a new link from $T$ to the percentage change in PPI which is not in any country-specific causal graphs; we interpret this as finding a new relationship through increasing the statistical power by pooling the data. Note that Country does not point to any other variable suggesting that the behavior across countries is independent of the country; this observation could be used to justify training a single econometric model across countries.
Figure 13: Results of CD-NOTS KCIT ran on the economic variables of multiple countries.
@@
Similar to the previous case study, we see statistically significant lagged relationships and statistically significant nonstationarity, implying we would not be able to trust the results of an algorithm that cannot handle these types of properties. For our linearity test, we test the linearity of the percentage change of CPI and each parent node (both including and excluding the contemporaneous percentage change in PPI since its direction was not inferred). We find that the only time the linearity hypothesis cannot be rejected is when using the contemporaneous percentage change in PPI as a parent and checking the linearity with the one-lag percentage change of CPI. We note that this is inconsistent with our usage of ParCorr in Figure 12 since ParCorr is a linear-based CI test. In Appendix C, we show the results of using RCoT where we find qualitatively that the graphs are similar.
@@ 4.3 Company Financials and Returns
(a) Contemporaneous Effects
(b) Lagged Effects
(c) Contemporaneous Effects
(d) Lagged Effects
Figure 14: Results of CD-NOTS KCIT run on financials data for log Price-to-Book ratio (LPTB) and stock returns (Ret). Normalized LPTB (N. LPTB) refers to normalizing the LPTB by the mean and standard deviation of the LPTB of financial companies in that quarter.
@@
For our third case study, we analyze the relationship between the Price-to-Book ratio and returns of financial companies in the S&P 500. We use data from the beginning of 2010 to the end of 2023. We utilize reported quarterly numbers, where we set the date to be the end of each quarter. To ensure no look-ahead bias when defining the Price-to-Book ratio, for any given date (i.e., end of quarter), we use the Price-to-Book ratio available that day. We join the next quarter returns with this company’s financial information. For our experiments, we use the log Price-to-Book ratio (LPTB) as well as the normalized LPTB, which we define as the LPTB normalized by the mean and standard deviation of the LPTB of financial companies in that quarter.
@@
In Figure 14(b) and Figure 14(d), we compare the graphs we get running CD-NOTS using KCIT using normalized LPTB instead of LPTB. We find that the nonstationarity is removed (T does not point to N. LPTB). Further, we see that the previous quarter’s normalized LPTB points to the next quarter’s returns with a reasonable p-value suggesting that its inclusion in the graph is not purely a function of the threshold we chose. In Figure 15(a), we compare the next quarter returns and the normalized LPTB and see that there is a small trend in the mean (the black circles). To this end, we create a simple trading strategy where we invest in the subset of the financial securities whose LPTB is greater than some threshold. In Figure 15(b), we see this simple approach outperforms a market-cap weighted strategy; prior work (Monge et al., 2023) has found that growth stocks have been outperforming value stocks. Importantly, we note that the goal of this is to not introduce a new signal but to show how causal discovery can be a tool in searching for statistically significant relations which, if the assumptions are true for the data, can be interpreted as causal.
(a) Comparison of Returns and N. LPTB
(b) Cumulative Returns of Strategies
Figure 15: Testing out predictive power of Normalized LPTB for returns of financial companies. For Figure 15(b), our strategy is long-ing financial companies that have a normalized LPTB greater than the threshold which we compare against the market-cap weighted portfolio.
@@
Similar to previous case studies, we see statistically significant lagged relationships and statistically significant nonstationarity, implying we would not be able to trust the results of an algorithm that cannot handle these types of properties. For our linearity test, we test the linearity of the normalized LPTB and returns conditional on all other parent nodes of returns and find the p-value to be 0.0 (i.e., we can reject the null hypothesis that the relationship is linear).